OpenStudy (anonymous):

Medal and fan help please Henry has 15 lawns mowed out of the total of 60 lawns.what percent of the lawns does Henry still have to mow show work plz!

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Let's try with simple division.

OpenStudy (anonymous):

How many times can 15 go into the number 60?

OpenStudy (anonymous):

Umm 10 not sure

OpenStudy (anonymous):

Multiply 15 a few times till you get to 60

OpenStudy (anonymous):

Or add 15 a few times by itself until you get to 60.

OpenStudy (anonymous):

15 times 2 equal 30 30 times 2 equal 60

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

An easier way of doing this would be by doing 15 times 3

OpenStudy (anonymous):

Which would get you 3

OpenStudy (anonymous):

Oh so help me fast plz

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

15+15+15+15=60. there are 4 15's in the number 60

OpenStudy (anonymous):

Since it takes 4 15's to make the number 60, that means the one 15 is a quarter of 60

OpenStudy (anonymous):

So what should I show work

OpenStudy (anonymous):

1/2=50%, while 1/3=33.3333367 percent, while 1/4=25%

OpenStudy (anonymous):

Plz tell me it clearly

OpenStudy (anonymous):



OpenStudy (anonymous):

Then

OpenStudy (anonymous):

Help me get the answer